# Manifesto from Dagstuhl Perspectives Workshop 24352—Conversational Agents: A Framework for Evaluation (CAFE)

### v1

@@ 7 Criteria
This section discusses the criteria to consider when instantiating an evaluation framework for the conversational search scenario. As mentioned in Section 4, a comprehensive evaluation of the cooperative interaction flow between a CONIAC system and its users should involve the goals of different stakeholders. In this sense, the evaluation criteria should also account for the different perspectives on the evaluation, e.g., from a user-oriented, company-focused, or even societal perspective [59, 116]. Sometimes, a stakeholder’s goals and the corresponding evaluation criteria can overlap and may account for multiple aspects. However, several evaluation criteria can be clearly attributed to the different aspects of the conversational search and recommendation process, being both objective and subjective.
Fig. 9 illustrates a taxonomy of criteria that can be used to evaluate a conversational system. The first partitioning of the criteria concerns the subject of the evaluation, and it allows us to identify two major classes of evaluation criteria: System-centric and User-centric evaluation criteria. By no means do we claim the listed criteria in the taxonomy to be exhaustive, but rather, we aim to categorize the different aspects of the cooperative conversational process, providing examples of suitable evaluation criteria.
@@
Hardware-centric evaluation criteria include, among others, the system’s energy usage, computational costs, efficiency, and latency. Modern state-of-the-art CONIAC systems mostly rely on LLMs with dedicated hardware and energy requirements. As earlier work by Strubell et al. [110] pointed out, the energy usage of LLMs has a considerable environmental impact, and these implications should be kept in mind by platform operators, system developers, and also by the users. Closely related to energy usage, computational costs are a more general criterion that is of special interest to the platform operators and service providers, who eventually have to cover the resulting financial costs. Regarding the system’s response time, latency and efficiency are important criteria for system operators and users alike. Long response times may dissatisfy the users, making the service less attractive, which can potentially lead to user abandonment or quitting the use of the service entirely.
@@
Other criteria—of particular interest to the society as a whole—include fairness, replicability, and the alignment to societal values. An evaluation of fairness criteria allows for a better understanding of potential biases in the data or the algorithms. Similarly, evaluating the replicability of system-based components allows audits of the reliability and robustness of the CONIAC system, which may inherently impact other criteria like the system’s trustworthiness. In this regard, the development of the system can also be driven by societal values, which, in turn, requires an evaluation of how well the CONIAC system is able to grasp and convey these values to the users in the end. In this context, veracity plays an important role, which accounts for the system’s ability to provide responses that also reflect reality. This is especially important, considering the consequences for decision-making contexts. Furthermore, recent work by Joko et al. [62] suggests that LLM-based conversational agents are not diverse, so diversity can be an important criterion for future evaluations of CONIAC systems.
We refer the reader to Sakai [97], who also provides an elaborated taxonomy of criteria for evaluating textual CONIAC systems. Many of the criteria mentioned by Sakai [97] can be aligned with the different aspects we envision in our taxonomy.
@@
User-centric criteria represent the second broader category of aspects that should be evaluated when it comes to interactive systems Kelly [65], Knijnenburg et al. [72] and especially CONIAC systems [61]. Users’ subjective reflections on their interaction with the system can be characterized as their perceptions of certain qualities of the system itself and the self-relevant consequences of their use of the system. For CONIAC systems, we divide users’ perceptions of system qualities into perceptions of the conversation itself (e.g., engagement, response quality), and perceptions of the content that is recommended/retrieved by the system (e.g., perceived accuracy, serendipity).
@@
It is important to note that the temporal dimension of the interaction is particularly important in CONIAC systems—this relates to both the flow of the conversation between the user and the system, the continuance of the cooperative interaction between the user and the system towards fulfilling the information access goal, and the development of the user’s trust in the system [64, 10]. This means that evaluating CONIAC systems requires us to evaluate the system at various temporal scopes: evaluation can address a single turn (i.e., the traditional evaluation metrics), a conversation (e.g., flow, continuance), or the system usage across multiple conversations (e.g., trust). Measuring the interaction at all of these levels means applying methodologies and metrics that extend beyond a single interaction and that do not just cover a CONIAC systems’ stationary properties but also its dynamic properties (which can be measured either continuously or at a certain discrete temporal rate).
@@ 8 Methodology
The advantage of CONIAC systems is their support for complex, rich, cooperative interactions that go beyond the back-and-forth interactions with traditional IR and RS systems. This richness of conversations must hence be reflected in the evaluation of CONIAC systems, even beyond extensions for generative systems like that of Gienapp et al. [46]. Traditional interactive IR and RS evaluations span from system-focused evaluations to user-focused evaluation methodologies [65, 129]—in the context of CONIAC systems, we advocate for an even stronger emphasis on user-focused evaluation to better capture the qualities of these complex, cooperative interactions. Additionally, the dimension of time is critical for CONIAC systems, which is not reflected in most IR and RS evaluation methodologies used so far.
@@
Evaluation methodologies for information systems can be placed along a continuum [65] from focusing on the system (without integrating actual users, like offline simulation) to focusing on the users (without integrating specific systems, like behavioral diary studies [115]). Methodologies focusing on the interaction between specific users and specific systems are placed in the center. The space depicted in Fig. 10 has this continuum as its horizontal axis.
Figure 10: In conversational settings, time makes the difference. See above the methodologies spectrum, “Involvement” (the x-axis) as introduced by Kelly [65] (shown lower left), extended by an orthogonal “Time model” axis. The figure also gives four examples of methodologies and how they cover different parts of the user involvement and time model space.
Conversations are complex, dynamic interactions that evolve continuously as they progress toward the user’s goal. This temporal aspect corresponds to two different concepts: The scope (how much we measure, from single turns to multi-conversation interactions) and the model of time (unit of analysis, from single discrete to continuous). We expand the one-dimensional continuum of Kelly [65] with the time model that the methodology implements, taken from the theory of system behavior [127] (cf. Fig. 10). Therefore, evaluation methodologies must be selected or adapted to fit the concept of time the research question demands. The four different time models, each implemented by several methodologies, are the following:
@@
A single data point is measured per conversation (named “memoryless” in system behavior theory). For some research questions, especially when comparing systems, it is sufficient to measure either only once for a conversation—typically at the end—or to take only an aggregated measurement like an average. A typical example of system-focused evaluations with this time model are single-interaction experiments, which include the TREC CAsT [33, 32, 34, 88] and TREC iKAT [3] competitions in which systems are asked to respond to a user utterance with conversation history, or the Netflix prize in which algorithms are asked to increase recommendation performance across a dataset [16]. A typical example of user-focused evaluations with this time model are controlled user studies with a post-conversation survey [70].
@@
The occurrence of discrete events is measured. Typically employed events are the utterances in conversations with chat-based agents and speaker changes for voice-based agents. Other suitable resolutions are single words uttered, topic switches, when the user has completed some sub-tasks and asks the agent for new instructions (e.g., when the agent assists in cooking), or when the recommended song is finished. A typical example of evaluation methodology is interaction-log analysis, where the events are reconstructed from the logs. Examples of available search interaction-logs include MS MARCO [86] Conversational Search [^4] An artificial log generated based on real search engine usage
@@
and CIRQL [119]. Detailed behavioral logs may be created by employing advanced user tracking mechanisms [125]. If no logs of real interactions are available, offline simulation can be employed to generate interactions, for example, using software like SimIIR [131], GenIRSim [68], and UserSimCRS [1].
One evaluation methodology that follows a discrete event time model that is especially relevant to CONIAC is discourse analysis, which entails parsing the conversation according to theories from the discourse and dialog literature as an intermediary step for the evaluation. The parsed representation of the conversation follows, for example, rhetorical structure theory [82], models of argumentation [109], or the conversational roles model [104]—the last was specifically developed for CONIAC and builds on top of the theory of speech acts [101]. The parsed representation can then be employed as a model of the conversation [67], for example, to detect anomalous conversations [121]. Available datasets include SCS [117] and MANtIS [93].
@@
Measurements are taken at discrete and regular points of time according to a selected sampling rate, ranging from a fraction of seconds to days. As conversations are typically analyzed in terms of the events described above, this time model is not as frequent in CONIAC experiments. However, especially for wearable or ubiquitous conversational agents like ones integrated into smart watches or glasses, experience sampling [91]—which uses the discrete time model by definition—could be an especially suitable methodology to gain insights into natural usage behavior. Discrete-time evaluations may also be used in longitudinal studies to track the evolution of users’ interactions with the system. Finally, in descriptive research, diary studies can be used to obtain user input at set time intervals, e.g., on a daily basis [23].
@@
Measurements are taken continuously. Though measurements taken with computers might not be genuinely continuous, methodologies that use this time model could use a very high sampling rate to effectively resemble continuous measurements and thus allow for calculating measurement derivatives. Physiological measurements are often taken continuously, for example, saccades and dwell times in eye tracking studies [13]. Continuous measurements do not need to be without interruption. When evaluating several interaction sessions with CONIAC systems, the measurement can pause.
@@
Offline simulation involves testing a system or model using pre-recorded data instead of real-time data [55]. Offline A/B testing compares two versions of a system by running simulations on historical data to see which performs better. These methods help in evaluating changes without affecting live users.
@@
User simulation refers to methods where user behaviors are simulated by software. Especially, LLM-based approaches allow for simulating users’ interaction behaviors for various tasks such as conversational search and recommendation [102, 132].
@@
Online A/B testing involves comparing two versions of a system to see which performs better [73]. Users are randomly shown either version A or version B, and their interactions are tracked to measure effectiveness and/or efficiency.
@@
Discourse analysis is a research field for the offline study of transcripts from written or spoken conversations. Discourse analysis can be done on various levels of abstraction, ranging from the very fine-grained conversation analysis family of methods [103] to methods based on, e.g., rhetorical structures, social processes, or systemic linguistics. In general, discourse analysis examines how linguistic and, in some cases, para-linguistic mechanisms and elements are used to achieve communicative goals in recorded interactions and how they make it possible for participants to make sense of a conversation [48].
@@
Data donation methodologies involve different ways people can share their data for research or analysis [130]. This can include directly providing data, allowing access to existing data, or contributing anonymized data collected by devices. These methods help researchers gather valuable information while respecting privacy and consent. They also allow the gathering of information that sits across multiple systems or authorities.
@@
The Wizard of Oz methodology involves creating a prototype where users interact with a system they believe to be automated but can be operated by a human behind the screen [31, 71]. This enables researchers to test and refine the system’s design and functionality before fully developing the technology. Alternatively, for conversational experiments, two humans could observe their interactions as a first step, referred to as observational studies [117]. It is a cost-effective, rapid deployment way to gather user feedback and improve the system based on real interactions.
@@
Controlled user studies involve observing and testing users in a structured environment to evaluate how they interact with a system [70]. Longitudinal studies extend this by following the same users (commonly referred to as panels) over a longer period to see how their interactions and experiences change over time. Commonly explicit feedback is gathered after the study. These methods help researchers understand both immediate and long-term effects.
@@
Qualitative user interview methodologies involve conducting in-depth conversations with users to understand their experiences, thoughts, and feelings. These interviews typically follow a methodology (e.g., grounded theory [26]). This approach helps researchers gain an understanding of motivations.
@@
For planning an evaluation, one usually starts from the goals of the evaluation (see Section 4) and then defines the criteria of interest (see Section 7). In the next step, one needs to choose an evaluation methodology covering these criteria. Table 2 shows the methodologies described above and lists the criteria categories of Section 7 for which they are most suitable. As the table shows, there is a mostly clear divide between system- and user-oriented methodologies [65].
@@ 9 Measures
There are tutorials and reviews of measures in IR [99], RS [49], and software [40]. As with all measures, it is important to ensure they are used with care [44].
@@
Users’ subjective reflection on their interaction with the system can be measured using multi-item measurement scales, many of which can be captured by existing user-centric evaluation frameworks [61, 72] or established scales for specific aspects (e.g., for various dimensions of trust [85]. It is important to adapt these scales to the context of the system (e.g., a scale for goal achievement should broadly refer to the goal of the specific system under evaluation, if known) or to develop new scales where needed—the interested reader can refer to DeVellis [36] for a primer on how to develop robust subjective measurement scales.
Behavioral measures that comprise users’ interactions with the system can be measured using conversation logs and compiled into measures that have been defined extensively in the area of A/B testing [73]. Objective outcomes that occur outside the system have been covered in the domain of behavioral psychology [126].
It is important to note that users’ perceptions and behaviors (both within and outside the system) are subject to be affected by users’ personal characteristics and contextual variables [20]. It is also important to measure users’ perceptions and behaviors at the appropriate temporal scale and/or frequency. For example, users’ decision-making can be traced temporally through expansive interaction logging [100], and the development of trust within an interaction session can be measured by repeatedly probing users with trust-related questions or by repeatedly observing trust-related behaviors [63, 64].

### v2

@@ 7 Criteria
This section discusses the criteria to consider when instantiating an evaluation framework for the conversational search scenario. As mentioned in Section 4, a comprehensive evaluation of the cooperative interaction flow between a CONIAC system and its users should involve the goals of different stakeholders. In this sense, the evaluation criteria should also account for the different perspectives on the evaluation, e.g., from a user-oriented, company-focused, or even societal perspective [jannach2020_mcnamara, bauer2019_multimethod_multistakeholder, torkamaan2024]. Sometimes, a stakeholder’s goals and the corresponding evaluation criteria can overlap and may account for multiple aspects. However, several evaluation criteria can be clearly attributed to the different aspects of the conversational search and recommendation process, being both objective and subjective.
Fig. 9 illustrates a taxonomy of criteria that can be used to evaluate a conversational system. The first partitioning of the criteria concerns the subject of the evaluation, and it allows us to identify two major classes of evaluation criteria: System-centric and User-centric evaluation criteria. Note that the literature about criteria is rich and diversified, lacking standardized classifications for them across different research areas. It is beyond the scope of this manifesto to propose such a commonly agreed classification. Therefore, by no means do we claim the listed criteria in the taxonomy to be exhaustive, but rather, we aim to categorize the different aspects of the cooperative conversational process, providing examples of suitable evaluation criteria.
@@
Hardware-centric evaluation criteria include, among others, the system’s energy usage, computational costs, efficiency, and latency. Modern state-of-the-art CONIAC systems mostly rely on LLMs with dedicated hardware and energy requirements. As earlier work by DBLP:conf/acl/StrubellGM19 pointed out, the energy usage of LLMs has a considerable environmental impact, and these implications should be kept in mind by platform operators, system developers, and also by the users. Closely related to energy usage, computational costs are a more general criterion that is of special interest to the platform operators and service providers, who eventually have to cover the resulting financial costs. Regarding the system’s response time, latency and efficiency are important criteria for system operators and users alike. Long response times may dissatisfy the users, making the service less attractive, which can potentially lead to user abandonment or quitting the use of the service entirely.
@@
Other criteria—of particular interest to the society as a whole—include fairness, replicability, and the alignment to societal values. An evaluation of fairness criteria allows for a better understanding of potential biases in the data or the algorithms. Similarly, evaluating the replicability of system-based components allows audits of the reliability and robustness of the CONIAC system, which may inherently impact other criteria like the system’s trustworthiness. In this regard, the development of the system can also be driven by societal values, which, in turn, requires an evaluation of how well the CONIAC system is able to grasp and convey these values to the users in the end. In this context, veracity plays an important role, which accounts for the system’s ability to provide responses that also reflect reality. This is especially important, considering the consequences for decision-making contexts. Furthermore, recent work by DBLP:conf/sigir/JokoCRV0H24 suggests that LLM-based conversational agents are not diverse, so diversity can be an important criterion for future evaluations of CONIAC systems.
We refer the reader to DBLP:journals/corr/abs-2305-08290, who also provides an elaborated taxonomy of criteria for evaluating textual CONIAC systems. Many of the criteria mentioned by DBLP:journals/corr/abs-2305-08290 can be aligned with the different aspects we envision in our taxonomy.
@@
User-centric criteria represent the second broader category of aspects that should be evaluated when it comes to interactive systems DBLP:journals/ftir/Kelly09, knijnenburg2012 and especially CONIAC systems [jin2024]. Users’ subjective reflections on their interaction with the system can be characterized as their perceptions of certain qualities of the system itself and the self-relevant consequences of their use of the system. For CONIAC systems, we divide users’ perceptions of system qualities into perceptions of the conversation itself (e.g., engagement, response quality), and perceptions of the content that is recommended/retrieved by the system (e.g., perceived accuracy, serendipity).
@@
It is important to note that the temporal dimension of the interaction is particularly important in CONIAC systems—this relates to both the flow of the conversation between the user and the system, the continuance of the cooperative interaction between the user and the system towards fulfilling the information access goal, and the development of the user’s trust in the system [kahr2024-2, DBLP:journals/sigir/AzzopardiCKMTRAACREHHKK24]. This means that evaluating CONIAC systems requires us to evaluate the system at various temporal scopes: evaluation can address a single turn (i.e., the traditional evaluation metrics), a conversation (e.g., flow, continuance), or the system usage across multiple conversations (e.g., trust). Measuring the interaction at all of these levels means applying methodologies and metrics that extend beyond a single interaction and that do not just cover a CONIAC systems’ stationary properties but also its dynamic properties (which can be measured either continuously or at a certain discrete temporal rate).
@@ 8 Methodology
The advantage of CONIAC systems is their support for complex, rich, cooperative interactions that go beyond the back-and-forth interactions with traditional IR and RS systems. This richness of conversations must hence be reflected in the evaluation of CONIAC systems, even beyond extensions for generative systems like that of gienapp:2024a. Traditional interactive IR and RS evaluations span from system-focused evaluations to user-focused evaluation methodologies [DBLP:journals/ftir/Kelly09, zangerle2022_fevr]—in the context of CONIAC systems, we advocate for an even stronger emphasis on user-focused evaluation to better capture the qualities of these complex, cooperative interactions. Additionally, the dimension of time is critical for CONIAC systems, which is not reflected in most IR and RS evaluation methodologies used so far.
@@
Evaluation methodologies for information systems can be placed along a continuum [DBLP:journals/ftir/Kelly09] from focusing on the system (without integrating actual users, like offline simulation) to focusing on the users (without integrating specific systems, like behavioral diary studies [teevan:2004]). Methodologies focusing on the interaction between specific users and specific systems are placed in the center. The space depicted in Fig. 10 has this continuum as its horizontal axis.
Figure 10: In conversational settings, time makes the difference. See above the methodologies spectrum, “Involvement” (the x-axis) as introduced by Kelly [DBLP:journals/ftir/Kelly09] (shown lower left), extended by an orthogonal “Time model” axis. The figure also gives four examples of methodologies and how they cover different parts of the user involvement and time model space.
Conversations are complex, dynamic interactions that evolve continuously as they progress toward the user’s goal. This temporal aspect corresponds to two different concepts: The scope (how much we measure, from single turns to multi-conversation interactions) and the model of time (unit of analysis, from single discrete to continuous). We expand the one-dimensional continuum of DBLP:journals/ftir/Kelly09 with the time model that the methodology implements, taken from the theory of system behavior [wymore1976systems] (cf. Fig. 10). Therefore, evaluation methodologies must be selected or adapted to fit the concept of time the research question demands. The four different time models, each implemented by several methodologies, are the following:
@@
A single data point is measured per conversation (named “memoryless” in system behavior theory). For some research questions, especially when comparing systems, it is sufficient to measure either only once for a conversation—typically at the end—or to take only an aggregated measurement like an average. A typical example of system-focused evaluations with this time model are single-interaction experiments, which include the TREC CAsT [dalton2020treccast2019conversational, DaltonEtAl2020, dalton:2021, owoicho:2022] and TREC iKAT [aliannejadi:2023] competitions in which systems are asked to respond to a user utterance with conversation history, or the Netflix prize in which algorithms are asked to increase recommendation performance across a dataset [bennett2007]. A typical example of user-focused evaluations with this time model are controlled user studies with a post-conversation survey [knijnenburg2015].
@@
The occurrence of discrete events is measured. Typically employed events are the utterances in conversations with chat-based agents and speaker changes for voice-based agents. Other suitable resolutions are single words uttered, topic switches, when the user has completed some sub-tasks and asks the agent for new instructions (e.g., when the agent assists in cooking), or when the recommended song is finished. A typical example of evaluation methodology is interaction-log analysis, where the events are reconstructed from the logs. Examples of available search interaction-logs include MS MARCO [nguyen:2016] Conversational Search [^4] An artificial log generated based on real search engine usage
@@
and CIRQL [trippas2024reevaluating]. Detailed behavioral logs may be created by employing advanced user tracking mechanisms [willemsen2011]. If no logs of real interactions are available, offline simulation can be employed to generate interactions, for example, using software like SimIIR [zerhoudi:2022], GenIRSim [kiesel:2024c], and UserSimCRS [DBLP:conf/wsdm/AfzaliDB023].
One evaluation methodology that follows a discrete event time model that is especially relevant to CONIAC is discourse analysis, which entails parsing the conversation according to theories from the discourse and dialog literature as an intermediary step for the evaluation. The parsed representation of the conversation follows, for example, rhetorical structure theory [mann:1988], models of argumentation [stede:2018], or the conversational roles model [sitter:1992]—the last was specifically developed for CONIAC and builds on top of the theory of speech acts [searle:1969]. The parsed representation can then be employed as a model of the conversation [KieselEtAl2021], for example, to detect anomalous conversations [vakulenko:2019]. Available datasets include SCS [trippas2020towards] and MANtIS [penha:2019].
@@
Measurements are taken at discrete and regular points of time according to a selected sampling rate, ranging from a fraction of seconds to days. As conversations are typically analyzed in terms of the events described above, this time model is not as frequent in CONIAC experiments. However, especially for wearable or ubiquitous conversational agents like ones integrated into smart watches or glasses, experience sampling [pejovic2016]—which uses the discrete time model by definition—could be an especially suitable methodology to gain insights into natural usage behavior. Discrete-time evaluations may also be used in longitudinal studies to track the evolution of users’ interactions with the system. Finally, in descriptive research, diary studies can be used to obtain user input at set time intervals, e.g., on a daily basis [carter2005].
@@
Measurements are taken continuously. Though measurements taken with computers might not be genuinely continuous, methodologies that use this time model could use a very high sampling rate to effectively resemble continuous measurements and thus allow for calculating measurement derivatives. Physiological measurements are often taken continuously, for example, saccades and dwell times in eye tracking studies [beckers:2010]. Continuous measurements do not need to be without interruption. When evaluating several interaction sessions with CONIAC systems, the measurement can pause.
@@
Offline simulation involves testing a system or model using pre-recorded data instead of real-time data [DBLP:conf/sigir/HsuMMPPLLPYSB24]. Offline A/B testing compares two versions of a system by running simulations on historical data to see which performs better. These methods help in evaluating changes without affecting live users.
@@
User simulation refers to methods where user behaviors are simulated by software. Especially, LLM-based approaches allow for simulating users’ interaction behaviors for various tasks such as conversational search and recommendation [sekulic2024, Zhang2024simulate].
@@
Online A/B testing involves comparing two versions of a system to see which performs better [kohavi2009]. Users are randomly shown either version A or version B, and their interactions are tracked to measure effectiveness and/or efficiency.
@@
Discourse analysis is a research field for the offline study of transcripts from written or spoken conversations. Discourse analysis can be done on various levels of abstraction, ranging from the very fine-grained conversation analysis family of methods [sidnell2012handbook] to methods based on, e.g., rhetorical structures, social processes, or systemic linguistics. In general, discourse analysis examines how linguistic and, in some cases, para-linguistic mechanisms and elements are used to achieve communicative goals in recorded interactions and how they make it possible for participants to make sense of a conversation [grosz1986attention].
@@
Data donation methodologies involve different ways people can share their data for research or analysis [DBLP:conf/chi/ZannettouNAGGRR24]. This can include directly providing data, allowing access to existing data, or contributing anonymized data collected by devices. These methods help researchers gather valuable information while respecting privacy and consent. They also allow the gathering of information that sits across multiple systems or authorities.
@@
The Wizard of Oz methodology involves creating a prototype where users interact with a system they believe to be automated but can be operated by a human behind the screen [dahlback1993, knijnenburg2016]. This enables researchers to test and refine the system’s design and functionality before fully developing the technology. Alternatively, for conversational experiments, two humans could observe their interactions as a first step, referred to as observational studies [trippas2020towards]. It is a cost-effective, rapid deployment way to gather user feedback and improve the system based on real interactions.
@@
Controlled user studies involve observing and testing users in a structured environment to evaluate how they interact with a system [knijnenburg2015]. Longitudinal studies extend this by following the same users (commonly referred to as panels) over a longer period to see how their interactions and experiences change over time. Commonly explicit feedback is gathered after the study. These methods help researchers understand both immediate and long-term effects.
@@
Qualitative user interview methodologies involve conducting in-depth conversations with users to understand their experiences, thoughts, and feelings. These interviews typically follow a methodology (e.g., grounded theory [charmaz2015grounded]). This approach helps researchers gain an understanding of motivations.
@@
For planning an evaluation, one usually starts from the goals of the evaluation (see Section 4) and then defines the criteria of interest (see Section 7). In the next step, one needs to choose an evaluation methodology covering these criteria. Table 2 shows the methodologies described above and lists the criteria categories of Section 7 for which they are most suitable. As the table shows, there is a mostly clear divide between system- and user-oriented methodologies [DBLP:journals/ftir/Kelly09].
@@ 9 Measures
There are tutorials and reviews of measures in IR [Sanderson2010], RS [Gunawardana2022], and software [FentonBieman2014]. As with all measures, it is important to ensure they are used with care [Fuhr2017].
@@
Users’ subjective reflection on their interaction with the system can be measured using multi-item measurement scales, many of which can be captured by existing user-centric evaluation frameworks [jin2024, knijnenburg2012] or established scales for specific aspects (e.g., for various dimensions of trust [mcknight2002]. It is important to adapt these scales to the context of the system (e.g., a scale for goal achievement should broadly refer to the goal of the specific system under evaluation, if known) or to develop new scales where needed—the interested reader can refer to devellis2011 for a primer on how to develop robust subjective measurement scales.
Behavioral measures that comprise users’ interactions with the system can be measured using conversation logs and compiled into measures that have been defined extensively in the area of A/B testing [kohavi2009]. Objective outcomes that occur outside the system have been covered in the domain of behavioral psychology [winkler1973].
It is important to note that users’ perceptions and behaviors (both within and outside the system) are subject to be affected by users’ personal characteristics and contextual variables [Cai2022trust]. It is also important to measure users’ perceptions and behaviors at the appropriate temporal scale and/or frequency. For example, users’ decision-making can be traced temporally through expansive interaction logging [schulte2017], and the development of trust within an interaction session can be measured by repeatedly probing users with trust-related questions or by repeatedly observing trust-related behaviors [kahr2024, kahr2024-2].